# ViP2-CLIP: Visual-Perception Prompting with Unified Alignment for Zero-Shot Anomaly Detection

### v1

@@ 5.1 Setup
We adopt standard metrics for evaluation: for image-level anomaly detection, we report the Area Under the Receiver Operating Characteristic curve (AUROC), Average Precision (AP), and maximum F1 score (F1); for pixel-level anomaly segmentation, we report AUROC, the Area Under the Per-Region Overlap curve (AUPRO), and F1.
@@
We use the publicly released CLIP (ViT-L/14@336px) as the frozen backbone. The learnable prompts contain 10 tokens, 3 of which interact explicitly with the global visual embedding. In UTPA, we retain the top 50 anomalous patches at each alignment layer as image-level descriptors and, following prior work Zhou et al. (2024); Cao et al. (2024), apply cross-modal alignment at layers 6, 12, 18, and 24. We fine-tune ViP2-CLIP on MVTec AD’s test split and evaluate zero-shot performance on all other datasets; for MVTec AD, we fine-tune on VisA’s test set. Dataset-level metrics are averaged across all subclasses. All experiments are implemented in PyTorch 2.6.0 on a single NVIDIA L20 (42 GB). Further implementation details are provided in Appendix B.
@@ 5.2 Main Results
Zero-Shot Anomaly Detection on Industrial Datasets
Table 1 compares ViP2-CLIP against five representative baselines on seven industrial defect benchmarks. Our model delivers competitive results and outperforms all baselines in most datasets. The vanilla CLIP struggles, as its pre-training emphasizes general object semantics rather than anomaly patterns. WinCLIP and APRIL-GAN improve detection through handcrafted prompts and local features tuning. AnomalyCLIP employs object-agnostic prompts but remains limited by static prompt design and shallow cross-modal fusion. AdaCLIP further refines prompts in both visual and textual space, which boosts F1 but neglects AUPRO, leading to an imbalanced ZSAD. In contrast, ViP2-CLIP yields consistent improvements across all metrics, achieving state-of-the-art performance. Our ViP-Prompt creates fine-grained normal and abnormal prompts that align tightly with object features; paired with UTPA’s unified alignment, it markedly boosts accuracy at both image-level detection and pixel-level localization.
Figure 4 compares our approach with four strong baselines, illustrating that ViP2-CLIP better captures defects with sharper boundaries, corroborating the quantitative improvements. We also provide computational efficiency analyses in Appendix C.
Zero-Shot Anomaly Detection on Medical Datasets
To assess cross-domain transferability, we benchmark ViP2-CLIP on eight public medical defect datasets. As shown in Table 2, ViP2-CLIP achieves consistently superior detection performance on most medical datasets, generally outperforming prior methods. Qualitative examples in Figure 4 further support this claim: our model precisely localizes melanoma lesions in dermoscopy images and colonic polyps in endoscopic frames, demonstrating robust adaptability to disparate pathological patterns.
@@
To quantify the benefit of visual conditioning, we compare ViP2-CLIP against ViP2-CLIP_re, which only uses static learnable prompts without visual cues. Figure 5 reports image-level and pixel-level F1 gains on the seven industrial datasets. ViP2-CLIP achieves significantly higher pixel-level F1 on all datasets and improves image-level F1 on five of them, confirming that injecting global and local visual context into prompts markedly strengthens cross-modal alignment, thus boosting superior ZSAD performance.
@@
Base (37.8, 11.5, 7.0) (74.1, 87.6, 87.2) (43.5, 14.6, 2.8) (60.2, 66.3, 74.5)
+G1 (89.5, 79.9, 38.6) (71.8, 86.2, 88.5) (95.2, 90.9, 31.6) (79.7, 82.8, 79.5)
+G2 (87.7, 83.8, 42.4) (88.9, 94.9, 91.6) (95.2, 91.6, 35.5) (83.0, 85.9, 81.4)
+G3 (90.5, 87.1, 43.1) (91.2, 96.0, 92.0) (95.4, 92.2, 33.6) (88.5, 90.4, 84.8)
Table 3: Module ablation.
5.3 Ablation Study
Module Ablation
We isolate the impact of three key components: ViP-ICA (G1), ViP-FGP (G2), and UTPA (G3). Table 3 shows that adding each module yields a clear, consistent gain, validating their standalone effectiveness. Specifically, ICA adaptively injects global visual context into prompt embeddings, while FGP incorporates fine-grained local cues, enabling the prompts to capture precise irregularities. Furthermore, UTPA’s unified patch-level alignment elevates both detection and localization performance, demonstrating its complementary value.
State-Adjective Robustness Ablation
To verify our prompt’s robustness, we replace the original ‘good/damaged’ adjectives with several alternative pairs of similar meaning and retrain the model on MVTec-AD and VisA. As Table 4 shows, detection performance is virtually unchanged under every substitution, indicating high robustness to specific adjective choices. This resilience stems from our learnable prompt tokens, which autonomously capture normal and anomalous semantics, obviating the need for handcrafted text templates.
Figure 6: Top-k component ablation.
@@ 5.3 Ablation Study
Top-k Ablation
To assess how the image-level representation is affected by the number of anomalous patches, we vary $K\in\{1,10,20,50,100,200\}$ and examine the AUROC performance in Figure 6. At $K=1$, the model is overly sensitive to local noise, yielding a low AUROC. As $K$ increases, performance steadily improves, demonstrating that aggregating multiple high-scoring anomalous patches enhances robustness. On MVTec AD, performance plateaus when $K$ is at 50 or 100; As for VisA, the optimum occurs at $K=50$. Beyond these points, additional patches introduce mostly normal regions, diluting the anomaly signals and reducing accuracy. Balanced across all metrics, we select the top 50 most anomalous patches as image-level descriptors.

### v2

@@ 5.1 Setup
We adopt standard metrics for evaluation: for image-level anomaly detection, we report the Area Under the Receiver Operating Characteristic curve (AUROC), Average Precision (AP), and maximum F1 (F1); for pixel-level anomaly segmentation, we report AUROC, the Area Under the Per-Region Overlap curve (AUPRO), and F1.
@@
We use the publicly released CLIP (ViT-L/14@336px) as the frozen backbone. The learnable prompts contain 10 tokens, 3 of which interact explicitly with the global visual embedding. In UTPA, we retain the top 50 anomalous patches at each alignment layer as image-level descriptors and, following prior work [33, 4], apply cross-modal alignment at layers 6, 12, 18, and 24. We fine-tune ViP2-CLIP on MVTec AD’s test split and evaluate ZSAD performance on all other datasets; for MVTec AD, we fine-tune on VisA’s test set. Dataset-level metrics are averaged across all subclasses. All experiments are implemented in PyTorch 2.6.0 on a single NVIDIA L20 (48 GB). Further implementation details are provided in Appendix B.
@@ 5.2 Main Results
Zero‐Shot Anomaly Detection on Industrial Datasets
Tab. 1 compares ViP2-CLIP against five representative baselines on seven industrial defect benchmarks. Our model delivers competitive results and outperforms all baselines in most datasets. The vanilla CLIP struggles, as its pre-training emphasizes general object semantics rather than anomaly patterns. WinCLIP and APRIL-GAN improve detection through handcrafted prompts and local features tuning. AnomalyCLIP employs object-agnostic prompts but remains limited by static prompt design and shallow cross-modal fusion. AdaCLIP further refines prompts in both visual and textual space, which boosts F1 but neglects AUPRO, leading to an imbalanced ZSAD. In contrast, ViP2-CLIP yields consistent improvements across all metrics, achieving state-of-the-art performance. Our ViP-Prompt creates fine-grained normal and abnormal prompts that align tightly with object features; paired with UTPA’s unified alignment, it markedly boosts accuracy at both image-level detection and pixel-level localization.
Fig. 4 compares the anomaly detection results of our approach with four strong baselines, showing that ViP2-CLIP more accurately captures defects with sharper boundaries, supporting the observed quantitative gains. We also provide failure cases analyses in Appendix C.
Zero‐Shot Anomaly Detection on Medical Datasets
To assess cross-domain transferability, we benchmark ViP2-CLIP on eight public medical defect datasets. As shown in Tab. 2, ViP2-CLIP achieves consistently superior detection performance on most medical datasets, generally outperforming prior methods. Qualitative examples in Fig. 4 further support this claim: our model precisely localizes melanoma lesions in dermoscopy images and colonic polyps in endoscopic frames, demonstrating robust adaptability to disparate pathological patterns.
@@
To quantify the benefit of visual conditioning, we compare ViP2-CLIP against ViP2-CLIP_re, which only uses static learnable prompts without visual cues. Fig. 5 reports image-level and pixel-level F1 gains on the seven industrial datasets. ViP2-CLIP achieves higher pixel‐level F1 on all datasets and improves image‐level F1 on five of them. The slight drops on BTAD and DAGM are attributed to significant domain shifts from MVTec AD: BTAD features real-world industrial environments with complex lighting conditions, while DAGM contains synthetic textures with small, repetitive anomalies. These differences likely cause shifts in global visual embeddings, limiting ICA’s transferability and leading to imbalanced detection performance. This underscores the importance of domain consistency in transfer learning, as large domain gaps can hinder generalization. Despite minor image-level drops on a few datasets, ViP-Prompt consistently achieves SOTA performance across diverse detection scenarios, demonstrating that incorporating both global and local visual cues into prompts enhances cross-modal alignment and promotes robust generalization.
Model Training Time (h) Inference Time (ms) GPU Cost (GB) Trainable Parameters (M)
AnomalyCLIP 1.02 90.72 ± 0.16 2.75 5.56
AdaCLIP 2.24 134.20 ± 0.39 3.17 10.67
ViP2-CLIP 0.54 48.69 ± 0.19 2.21 4.15
Table 3: Comparison of computational efficiency on VisA.
UTPA ViP-Prompt MVTec AD VisA
ICA FGP Pixel-level Image-level Pixel-level Image-level
(37.8, 11.5, 7.0) (74.1, 87.6, 87.2) (43.5, 14.6, 2.8) (60.2, 66.3, 74.5)
(89.5, 79.9, 38.6) (71.8, 86.2, 88.5) (95.2, 90.9, 31.6) (79.7, 82.8, 79.5)
(90.0, 85.5, 42.7) (66.1, 82.6, 86.6) (95.3, 91.5, 33.4) (60.1, 66.4, 73.2)
(87.7, 83.8, 42.4) (88.9, 94.9, 91.6) (95.2, 91.6, 35.5) (83.0, 85.9, 81.4)
(89.8, 84.4, 40.2) (89.2, 95.0, 91.4) (94.9, 91.4, 31.3) (86.3, 89.1, 83.8)
(88.8, 82.1, 38.7) (85.2, 93.3, 89.0) (95.3, 91.9, 31.4) (87.6, 89.4, 83.8)
(90.5, 87.1, 43.1) (91.2, 96.0, 92.0) (95.4, 92.2, 33.6) (88.5, 90.4, 84.8)
Table 4: Module ablation.
Computational Efficiency Analysis
Tab. 3 compares the computational efficiency of the top three baselines. AnomalyCLIP modifies self-attention weights via DPAM, while AdaCLIP introduces additional learnable tokens and a multi-scale clustering module, both of which increase computational overhead. In contrast, ViP2‑CLIP employs only two lightweight adapters outside the frozen CLIP backbone, achieving the fastest training (0.54 h), inference (48.7 ms/image), and the lowest GPU memory usage (2.21 GB), making it well-suited for deployment in resource-constrained industrial environments.
5.3 Ablation Study & Discussion
Module Ablation
We isolate the contributions of three key components: ViP-ICA, ViP-FGP, and UTPA. Tab. 4 shows that adding each module yields a clear, consistent gain, validating their standalone effectiveness. Specifically, ICA adaptively injects global visual context into prompt embeddings, while FGP incorporates fine‐grained local cues, enabling the prompts to capture precise irregularities. Furthermore, UTPA consistently improves both detection and localization performance upon integration, demonstrating the general effectiveness of the unified patch-level alignment strategy in guiding prompt learning toward more discriminative representations.
Figure 6: Top-k component ablation.
@@ 5.3 Ablation Study & Discussion
CLIP_DUAL (69.6, 24.6, 12.0) (88.6, 94.9, 91.3) (93.3, 82.4, 24.2) (79.3, 82.4, 79.0)
CLIP_UTPA (89.6, 83.7, 37.4) (85.0, 93.3, 88.8) (94.4, 88.4, 26.2) (83.8, 86.7, 81.4)
Table 5: Ablation on the effectiveness of UTPA.
UTPA for Resolving Optimization Conflicts
We conduct an ablation study using a frozen CLIP backbone with learnable prompts for normal and anomalous classes to evaluate the effectiveness of UTPA in guiding network learning. The baseline (CLIP_DUAL) adopts a dual-branch strategy that aligns prompts with both global and local visual embeddings. In contrast, CLIP_UTPA simplifies the optimization by aligning solely with local features. As shown in Tab. 5, on the MVTec AD dataset, the dual-branch strategy tends to overfit image-level scores while underperforming on pixel-level detection. UTPA, by comparison, achieves more stable and superior results. As for VisA, it outperforms the baseline across both pixel- and image-level metrics. Several results in Tab. 4 also support this observation: traditional dual-branch approaches often suffer from imbalanced performance due to conflicting objectives, whereas UTPA alleviates this issue through consistent text-patch alignment. This facilitates more robust model learning and provides a novel alignment paradigm for more accurate ZSAD research.
Top‐K Ablation
To assess how the image-level representation is affected by the number of anomalous patches, we vary $K\in\{1,10,20,50,100,200\}$ and examine the AUROC performance in Fig. 6. At $K=1$, the model is overly sensitive to local noise, yielding a low AUROC. As $K$ increases, performance steadily improves, demonstrating that aggregating multiple high-scoring anomalous patches enhances robustness. On MVTec AD, performance plateaus when $K$ is at 50 or 100; As for VisA, the optimum occurs at $K=50$. Beyond these points, additional patches introduce mostly normal regions, diluting the anomaly signals and reducing accuracy. Balanced across all metrics, we select the top 50 most anomalous patches as image-level descriptors, which aligns well with the anomaly distribution characteristics in most detection scenarios.
Feature Visualization at Different layers
To investigate the role of network depth in capturing anomalous features, Fig. 7 presents anomaly maps derived from $F_{m}^{1}$ to $F_{m}^{4}$ layers. We observe that for complex object categories such as hazelnut, deeper layers yield more accurate anomaly localization. In contrast, for simpler texture-based objects like leather, intermediate layers offer sufficient discriminative power. Aggregating features across multiple depths provides complementary contextual cues, allowing the model to better adapt to diverse anomaly types, and thereby achieve more robust anomaly detection.
State‐Adjective Ablation
To verify our prompt’s robustness, we replace the original ‘good/damaged’ adjectives with several alternative pairs of similar meaning and retrain the model on MVTec-AD and VisA. As Tab. 6 shows, detection performance remains consistent across all variants, demonstrating insensitivity to specific adjective choices. This resilience stems from our learnable prompt tokens, which autonomously capture normal and anomalous semantics, obviating the need for handcrafted text templates.
Figure 7: Visualization of anomaly maps from different layers.